# VariFace: Fair and Diverse Synthetic Dataset Generation for Face Recognition

### v1

@@ 7 Face Vendi Score Guidance
8 Divergence Score Conditioning
8.1 Label noise detection with Divergence scores
@@
8.2 Hyperparameter evaluation
To demonstrate the effect of hyperparameter settings for the DS, we evaluate a broad range of DS conditioning values (Table. S2).
@@ 9 Age Conditioning
10 Synthetic faces comparison with real faces
@@
11 Diffusion model settings
The settings for the diffusion models used in VariFace are shown in Table S3. We generally follow the default settings from the original HDiT, except for using Adaptive Discriminator Augmentation. [26], and modifying the conditioning module to accommodate for multiple conditions (Fig. S7).
Figure S7: VariFace HDiT block. We follow the structure of the HDiT block used in [12] and modify the conditioning block to handle multiple conditions. The structure of the conditioning block is shown for the stage 2 model. For the stage 1 model, the identity condition is removed, and age and divergence scores are replaced with race and gender labels. GEGLU is used as the MLP [62].
@@ 11 Diffusion model settings
12 Face recognition model settings
The settings used for the FR model are shown in Table S4. We generally follow default settings from the InsightFace library [14], except for substituting the polynomial learning rate scheduler for a step learning rate scheduler and including additional data augmentations.
@@ 12 Face recognition model settings
13 Baseline real dataset performance
@@
Kim et al. 2024 [31] 0.9935 0.9597 0.8412 0.9365 0.9078 0.9277
Kim et al. 2023 [30]; Papantoniou et al. 2024 [49] 0.9942 0.9656 0.8973 0.9408 0.9332 0.9462
Boutros et al. 2023 [9]; Kolf et al. 2023 [32] 0.9955 0.9531 0.8995 0.9455 0.9378 0.9463
Wu et al. 2024 [72] 0.9938 0.9691 0.8978 0.9450 0.9335 0.9479
Boutros et al. 2023 [8] 0.9952 0.9552 0.9038 0.9477 0.9393 0.9482
@@
14 Effect of loss function on model performance
@@
ArcFace [14] 0.9950 0.9536 0.9007 0.9487 0.9368 0.9470 0.8822 0.8697 0.9448 0.8978 0.8986
CurricularFace [23] 0.9920 0.9421 0.8813 0.9308 0.9247 0.9342 0.8463 0.8340 0.9163 0.8608 0.8644
SphereFace [39] 0.9943 0.9524 0.9018 0.9478 0.9380 0.9469 0.8787 0.8658 0.9415 0.9003 0.8966
AdaFace [29] 0.9955 0.9534 0.9007 0.9485 0.9345 0.9465 0.8787 0.8648 0.9452 0.8950 0.8959
MagFace [45] 0.9940 0.9517 0.8952 0.9488 0.9332 0.9446 0.8800 0.8628 0.9453 0.8900 0.8945
UniFace [78] 0.9945 0.9531 0.9008 0.9437 0.9348 0.9454 0.8712 0.8583 0.9437 0.8917 0.8912
@@
Since the development of the ArcFace loss, there have been numerous alternative loss functions proposed in the literature [39, 23, 45, 29, 78]. The performance using different loss functions for FR models trained on the CASIA-WebFace dataset is shown in Table S6. Generally, the performance is similar across loss functions, with the best performance obtained using the ArcFace loss.
@@
SynFace [53] 0.5M 0.8580 0.6473 0.6395 0.5765 0.6987 0.6840 -0.2630 -
SynFace [53] 1.0M 0.8593 0.6341 0.6452 0.5750 0.6860 0.6799 -0.2671 -
DigiFace [2] 0.5M 0.8508 0.7431 0.6498 0.6093 0.6713 0.7049 -0.2421 -0.1296
DigiFace [2] 1.2M 0.8872 0.7827 0.6835 0.6218 0.7170 0.7384 -0.2086 -0.1148
DCFace [30] 0.5M 0.9863 0.8896 0.8325 0.9082 0.9173 0.9068 -0.0402 +0.0112
DCFace [30] 1.2M 0.9895 0.886 0.8497 0.9173 0.9268 0.9139 -0.0331 +0.0018
Vec2Face [72] 0.5M 0.9848 0.8737 0.8413 0.9187 0.9298 0.9097 -0.0373 -0.0103
Vec2Face [72] 1.0M 0.9868 0.8819 0.8537 0.9408 0.9372 0.9201 -0.0269 -0.0046
@@
15 Reproducing performance from open-source synthetic datasets
@@
Except for DigiFace, we observed a similar performance with our training settings and the original paper results. Our performance obtained with DigiFace is significantly lower than the results in the original paper, and this is likely due to differences in the data augmentation strategies used [2]. While our data augmentation settings are tuned for real data, DigiFace employs additional data augmentation to overcome the domain gap between real and CG data. Importantly, regardless of whether the original dataset performance or our reproduced performances are used, our proposed method remains the best-performing method.
16 Limitations

### v2

@@ Appendix B Face Vendi Score Guidance
Appendix C Divergence Score Conditioning
C.1 Label noise detection with Divergence scores
@@
C.2 Hyperparameter evaluation
To demonstrate the effect of hyperparameter settings for the DS, we evaluate a broad range of DS conditioning values (Table S2).
@@ Appendix D Age Conditioning
Appendix E Synthetic faces comparison with real faces
@@
Appendix F Diffusion model settings
The settings for the diffusion models used in VariFace are shown in Table S3. We generally follow the default settings from the original HDiT, except for using Adaptive Discriminator Augmentation. (Karras et al., 2020), and modifying the conditioning module to accommodate for multiple conditions (Fig. S7).
Figure S7: VariFace HDiT block. We follow the structure of the HDiT block used in (Crowson et al., 2024) and modify the conditioning block to handle multiple conditions. The structure of the conditioning block is shown for the stage 2 model. For the stage 1 model, the identity condition is removed, and age and divergence scores are replaced with race and gender labels. GEGLU is used as the MLP (Shazeer, 2020).
@@ Appendix F Diffusion model settings
Appendix G Face recognition model settings
The settings used for the FR model are shown in Table S4. We generally follow default settings from the InsightFace library (Deng et al., 2019), except for substituting the polynomial learning rate scheduler for a step learning rate scheduler and including additional data augmentations.
@@ Appendix G Face recognition model settings
Appendix H Baseline real dataset performance
@@
Kim et al. 2024 (Kim et al., 2024) 0.9935 0.9597 0.8412 0.9365 0.9078 0.9277
Kim et al. 2023 (Kim et al., 2023); Papantoniou et al. 2024 (Papantoniou et al., 2024) 0.9942 0.9656 0.8973 0.9408 0.9332 0.9462
Boutros et al. 2023 (Boutros et al., 2023b); Kolf et al. 2023 (Kolf et al., 2023) 0.9955 0.9531 0.8995 0.9455 0.9378 0.9463
Wu et al. 2024 (Wu et al., 2024a) 0.9938 0.9691 0.8978 0.9450 0.9335 0.9479
Boutros et al. 2023 (Boutros et al., 2023a) 0.9952 0.9552 0.9038 0.9477 0.9393 0.9482
@@
Appendix I Effect of loss function on model performance
@@
ArcFace (Deng et al., 2019) 0.9950 0.9536 0.9007 0.9487 0.9368 0.9470 0.8822 0.8697 0.9448 0.8978 0.8986
CurricularFace (Huang et al., 2020) 0.9920 0.9421 0.8813 0.9308 0.9247 0.9342 0.8463 0.8340 0.9163 0.8608 0.8644
SphereFace (Liu et al., 2017) 0.9943 0.9524 0.9018 0.9478 0.9380 0.9469 0.8787 0.8658 0.9415 0.9003 0.8966
AdaFace (Kim et al., 2022b) 0.9955 0.9534 0.9007 0.9485 0.9345 0.9465 0.8787 0.8648 0.9452 0.8950 0.8959
MagFace (Meng et al., 2021) 0.9940 0.9517 0.8952 0.9488 0.9332 0.9446 0.8800 0.8628 0.9453 0.8900 0.8945
UniFace (Zhou et al., 2023) 0.9945 0.9531 0.9008 0.9437 0.9348 0.9454 0.8712 0.8583 0.9437 0.8917 0.8912
@@
Since the development of the ArcFace loss, there have been numerous alternative loss functions proposed in the literature (Liu et al., 2017; Huang et al., 2020; Meng et al., 2021; Kim et al., 2022b; Zhou et al., 2023). The performance using different loss functions for FR models trained on the CASIA-WebFace dataset is shown in Table S6. Generally, the performance is similar across loss functions, with the best performance obtained using the ArcFace loss.
@@
SynFace (Qiu et al., 2021) 0.5M 0.8580 0.6473 0.6395 0.5765 0.6987 0.6840 -0.2630 -
SynFace (Qiu et al., 2021) 1.0M 0.8593 0.6341 0.6452 0.5750 0.6860 0.6799 -0.2671 -
DigiFace (Bae et al., 2023) 0.5M 0.8508 0.7431 0.6498 0.6093 0.6713 0.7049 -0.2421 -0.1296
DigiFace (Bae et al., 2023) 1.2M 0.8872 0.7827 0.6835 0.6218 0.7170 0.7384 -0.2086 -0.1148
DCFace (Kim et al., 2023) 0.5M 0.9863 0.8896 0.8325 0.9082 0.9173 0.9068 -0.0402 +0.0112
DCFace (Kim et al., 2023) 1.2M 0.9895 0.886 0.8497 0.9173 0.9268 0.9139 -0.0331 +0.0018
Vec2Face (Wu et al., 2024a) 0.5M 0.9848 0.8737 0.8413 0.9187 0.9298 0.9097 -0.0373 -0.0103
Vec2Face (Wu et al., 2024a) 1.0M 0.9868 0.8819 0.8537 0.9408 0.9372 0.9201 -0.0269 -0.0046
@@
Appendix J Reproducing performance from open-source synthetic datasets
@@
Except for DigiFace, we observed a similar performance with our training settings and the original paper results. Our performance obtained with DigiFace is significantly lower than the results in the original paper, and this is likely due to differences in the data augmentation strategies used (Bae et al., 2023). While our data augmentation settings are tuned for real data, DigiFace employs additional data augmentation to overcome the domain gap between real and CG data. Importantly, regardless of whether the original dataset performance or our reproduced performances are used, our proposed method remains the best-performing method.
Appendix K Limitations